The area of a trangle is 13.5 square meters. Find the base and height of the rectangle if its height is 6 meters greater than it base. Use an equation and the formula: area of a triangle= 0.5(base)(height): The area of a trangle is 13.5 square meters. Find the base and height of the rectangle if its height is 6 meters greater than it base. Use an equation and the formula: area of a triangle= 0.5(base)(height)
Answer by checkley77(3624)

H=B+6
A=BH/2
13.5=B(B+6)/2
13.5=(B^2+6B)/2
2*13.5=B^2+6B
B^2+6B-27=0
(B+9)(B-3)=0
B-3=0
B=3 FOR THE BASE
3+6=9 FOR THE HEIGHT.
PROOF:
13.5=(3*9)/2
13.5=27/2
13.5=13.5